OpenStudy (axel.caballero):

how is adding and subtracting rational expressions similar to adding and subtracting simple fraction ?

OpenStudy (ranga):

In both cases you make the denominator common and then add or subtract the numerator. To add fractions 1/2 and 1/3, the common denominator is 6 1/2 = (1*3) / (2*3) = 3/6 1/3 = (1*2) / (3*2) = 2/6 Add 1/2 + 1/3 = 3/6 + 2/6 = (3+2) / 6 = 5/6 You do the same with rational expressions. Add: 1/x + 1/(x + 1). Common denominator is x(x+1) 1/x + 1/(x + 1) = (x+1) / {x(x+1)} + x / {x(x+1)} = (x+1+x) / {x(x+1)} = (2x+1)/{x(x+1)}